40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 21 18:18:42 2022, yehonathan, Update, BHD, POX Alignment POX11_alignment.png
    Reply  Thu Apr 28 17:45:53 2022, Anchal, Update, BHD, POX Beam issues PXL_20220429_002703686.jpg
       Reply  Sat Apr 30 13:29:33 2022, Paco, Update, BHD, POX Beam issues --- fixed, ASL alignment on vertex chamber 
          Reply  Mon May 2 17:51:30 2022, Anchal, Update, BHD, POX Beam aligned on RFPD, YARM locked 
Message ID: 16824     Entry time: Mon May 2 17:51:30 2022     In reply to: 16821
Author: Anchal 
Type: Update 
Category: BHD 
Subject: POX Beam aligned on RFPD, YARM locked 

[Paco, Anchal]

We found that one of the Y1-1037-45P marked mirror that we used was actually curved. So we removed it and used a different Y1-1037-45P mirror, adjusted the position of the lens and got the beam to land on POX11 RFPD successfully.

Then in control room, we maximized the POX11_I_ERR PDH signal amplitude by changing C1:LSC-POX11_PHASE_R to 42.95 from -67.7. We kept the C1:LSC-POX11_PHASE_D same at 90. We were getting +/- 200 PDH signal on POX_I_ERR.

Then in our attempt to lock the XARM, when we ran the "Restore XARM (POX)" script, YARM locked!

We are not sure why the YARM locked, we might have gotten lucky today. So we ran ASS on YARM and got the transmission (TRY_OUT) stable at 1. The lock is very robust and retrievable.

Coming back to XARM, we realized that the transmission photodiode used for XARM was the low-gain QPD instead of the thorlabs high gain photodiode. The high-gain photodiode was outputing large negative counts for some reason. We went to the Xend to investigate and found that the high gain photodiode was disconnected for some reason. Does anyone know/remember why we disconnected this photodiode?

We connected the photodiode back and it seems to work normally. We changed the photodiode selection back to high gain photodiode for TRX and on 40 dB attenuation, we see flashing between 1.4 to 1.6. However, we were unable to lock the XARM. We tried changing the gain of the loop, played a little bit with the trigger levels etc but couldn't get it to lock. Next shift team, please try to lock XARM.

Quote:

[Paco, Anchal, Yuta]

We opened the BSC and ITMX chamber in the morning (Friday) to investigate POX11 beam clipping. We immediately found that the POX11 beam was clipping by the recently installed cable posts, so luckily no major realingment had to be done after reinstalling the cable post in a better location.


Because we had the BSC open, we decided to steer the AS1 mirror to align the AS path from ITMY all the way to the vertex chamber.  Relatively small AS1 offsets (of ~ 2000 counts each) were added on PIT / YAW to center the beam on ASL (there is slight clipping along PIT, potentially because of the AS2 aperture. We then opened the vertex chamber and located the AS beam with relative ease. We decided to work on this chamber, since major changes propagate heavily downstream (simply changing the IMC pointing).

Anchal removed old optics from the vertex chamber and we installed the steering pair of mirrors for AS path. This changed the balance of the vertex table by a lot. By using the MC REFL camera beam spot we managed to coarsely balance the counterweights and recover the nominal IMC injection pointing. Simply reenabling the IMC autolocker gave us high transmission (~ 970 counts out of the typical 1200 these days).

The final IMC alignment was done by Anchal with delicate PIT motion on the input injection IMC miror to maximize the transmission (to our satisfaction, Anchal's motion was fine enough to keep the IMC locked). The end result was quite satisfying, as we recovered ~ 1200 counts of MC transmission.

Finally, we looked at the arm cavity transmission to see if we were lucky enough to see flashing. After not seeing it, we adjusted TT1 / TT2 to correct for any MMTT1 pitch adjustment needed after the vertex table rebalancing. Suprisingly, we didn't take too long and recovered the nominal arm cavity pointing after a little adjustment. We stopped here, but now the vertex table layout is final, and AS beam still needs to be aligned to the vertex in-air table.

 

ELOG V3.1.3-